OpenStudy (anonymous):

when ever u r doing decimals is .8 higher than .27

OpenStudy (shubhamsrg):

convert them into fractions if you are confused 0.8 = 8/10 .27=27/100 try to make denominators same i.e. 8/10 = 80/100 now compare both. hope it helps ?